OpenStudy (anonymous):

A gardener purchases a ceramic planter, in the shape of a hemisphere, for a small batch of leftover annuals. The volume of a hemisphere is modeled by the function V = 2/3 πr 3 Write a model for the radius as a function of the volume. The label on the planter says that it holds approximately 134 cubic inches of potting soil. What is the radius of the planter, rounded to the nearest inch? Use 3.14 for π

OpenStudy (anonymous):

You would need to invert your equation. IE solve for r. So we need everything at the end to equal r. Do you need help?

OpenStudy (anonymous):

Yes, major help, I am new to this.

OpenStudy (anonymous):

OK our equation is V=2/3 (pi)(r^3) Correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

I'm sorry the right side of the equation?

OpenStudy (anonymous):

right side of the equation???

OpenStudy (anonymous):

Never mind we need to reflect the answer with pi.

OpenStudy (anonymous):

V=2/3 (pi)(r^3) So we need to move things from the right side to the left. To get rid of 2/3 from the right side what would we have to do to both sides of the equation?

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

BY what?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

3 times 2/3 = 6/3 = 2. We need to get to 1. So 3 is incorrect. What would you multiply by 2/3 to get 1?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

Correct so if we multiply both sides by 3/2 waht will our equation look like?

OpenStudy (anonymous):

pi(r^3)

OpenStudy (anonymous):

Actually \[\frac{ 3}{ 2 }V= \pi (r ^{3})\] What ever we do to one side we must do to the other. Now how do we make (pi) equal 1?

OpenStudy (anonymous):

multiply 3/2 x 2/3

OpenStudy (anonymous):

How do you make (pi) = 1. If 1 = pi/pi.

OpenStudy (anonymous):

I'm guessing divide pi by pi?

OpenStudy (anonymous):

Any value multiplied by its reciprocal will equal 1. Examble attached\[2 \times \frac{ 1 }{ 2 }=1.....or.....3\times \frac{ 1 }{ 3 }=1\] So (pi) x what = 1

OpenStudy (anonymous):

(pi) times what will equal 1?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

The reciprocal of (pi) is 1/(pi)...\[\frac{ \pi }{ 1 }\times \frac{ 1 }{ \pi }= \frac{ \pi }{ \pi }..which equals..1\] Do you see this?

OpenStudy (anonymous):

(pi) over one is the same as (pi)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what does frac means

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Now what does the equation look like when you multiply both sides by (pi)/1. Remember what ever you do to one side you must do to the other.

OpenStudy (anonymous):

''[\frac{ \pi }{ 1 }\times \frac''

OpenStudy (anonymous):

It was the equation (pi)/1 times 1/(pi) =1 but it must not have converted on your screen

OpenStudy (anonymous):

ok cause how it was written out I never seen it written like that before

OpenStudy (anonymous):

Now what does the equation look like when you multiply both sides by (pi)/1. Remember what ever you do to one side you must do to the other.